# Unlocking the Potential of Model Calibration in Federated Learning

### v2

@@ B.2 Additional Results of Calibration on FL
Calibration Method FedAvg McMahan et al. [2017] FedProx Sahu et al. [2018] Scaffold Karimireddy et al. [2019] FedDyn Acar et al. [2021] FedNova Wang et al. [2020b]
Uncal. 97.19 0.64 0.26 97.50 0.92 0.30 97.59 0.69 0.23 99.04 0.93 0.39 97.92 0.74 0.27
Focal Lin et al. [2017] 97.70 3.84 0.83 97.35 4.70 0.99 97.45 2.76 0.62 99.01 3.48 0.58 97.76 3.92 0.82
LS Müller et al. [2019] 97.97 13.42 2.71 97.72 13.79 2.77 97.87 13.22 2.67 94.38 4.38 0.52 97.98 13.39 2.70
BS Brier [1950] 95.93 2.72 0.64 95.24 3.03 0.71 96.29 2.09 0.51 91.72 8.28 0.83 96.39 2.92 0.63
MMCE Kumar et al. [2018] 98.48 5.40 1.12 98.19 5.51 1.14 95.10 15.22 0.68 98.08 1.51 0.29 98.49 5.41 1.12
FLSD Mukhoti et al. [2020] 97.23 11.74 2.33 96.68 13.21 2.62 97.12 9.66 1.91 97.28 1.48 0.39 97.20 11.95 2.37
MbLS Liu et al. [2021] 98.01 2.93 0.66 97.75 3.39 0.74 97.91 2.65 0.61 98.32 2.39 0.49 98.07 2.97 0.67
DCA Liang et al. [2020] 98.13 0.31 0.19 97.83 0.36 0.22 97.71 0.51 0.21 98.93 0.57 0.23 98.11 0.26 0.19
MDCA Hebbalaguppe et al. [2022b] 97.87 0.60 0.25 97.58 0.91 0.30 97.68 0.48 0.23 99.01 0.61 0.24 97.81 0.63 0.24
NUCFL (DCA+COS) 98.97 0.56 0.15 98.50 0.42 0.18 97.40 0.33 0.20 98.88 0.54 0.27 99.02 0.24 0.16
NUCFL (MDCA+COS) 98.94 0.54 0.15 98.55 0.27 0.15 96.97 0.35 0.21 98.83 0.53 0.21 98.99 0.49 0.15
NUCFL (DCA+L-CKA) 98.99 0.59 0.16 98.77 0.41 0.14 96.84 0.32 0.28 98.63 0.33 0.16 98.08 0.25 0.15
NUCFL (MDCA+L-CKA) 98.99 0.57 0.15 98.89 0.31 0.16 97.11 0.32 0.28 98.69 0.31 0.15 98.99 0.23 0.15
NUCFL (DCA+RBF-CKA) 99.01 0.60 0.15 98.91 0.42 0.16 96.93 0.33 0.27 98.75 0.38 0.19 98.96 0.29 0.16
NUCFL (MDCA+RBF-CKA) 99.09 0.60 0.16 98.77 0.37 0.15 97.54 0.31 0.23 98.86 0.32 0.15 98.28 0.26 0.16
@@
Table 7: Accuracy (%), calibration measures ECE (%), and SCE (%) of various federated optimization methods with different calibration methods under IID scenario on the MNIST dataset. Values in boldface indicate the best calibration across all methods, achieved without compromising accuracy compared to the uncalibrated model.
@@
Calibration Method FedAvg McMahan et al. [2017] FedProx Sahu et al. [2018] Scaffold Karimireddy et al. [2019] FedDyn Acar et al. [2021] FedNova Wang et al. [2020b]
Uncal. 96.52 1.59 0.60 96.05 1.74 0.60 95.65 6.24 0.96 98.37 1.48 0.38 96.55 1.24 0.49
Focal Lin et al. [2017] 95.88 5.96 1.44 94.94 6.67 1.58 95.01 5.15 1.36 98.28 1.71 0.26 96.06 5.88 1.33
LS Müller et al. [2019] 97.53 20.50 4.13 96.65 2.41 4.14 96.75 2.16 4.40 98.80 5.15 3.06 97.48 1.91 3.84
BS Brier [1950] 94.26 3.52 1.07 93.58 4.28 1.06 94.27 3.94 1.06 94.68 5.31 0.53 93.96 3.15 0.88
MMCE Kumar et al. [2018] 98.43 7.32 1.51 94.74 9.06 1.85 86.48 15.38 4.52 97.48 1.58 0.49 98.30 7.25 1.47
FLSD Mukhoti et al. [2020] 93.95 14.96 3.38 93.29 16.31 3.52 93.41 13.88 3.13 95.84 1.99 0.39 94.93 15.07 3.15
MbLS Liu et al. [2021] 97.22 5.49 1.23 96.46 6.34 1.40 96.55 5.70 1.31 98.71 2.45 0.55 97.24 5.17 1.12
DCA Liang et al. [2020] 96.94 0.64 0.44 96.04 0.77 0.51 96.42 0.45 0.60 98.47 1.14 0.22 97.13 0.64 0.38
MDCA Hebbalaguppe et al. [2022b] 96.92 1.48 0.56 96.02 1.63 0.63 95.82 0.96 0.61 98.93 1.13 0.21 96.98 1.09 0.43
NUCFL (DCA+COS) 98.86 0.51 0.19 98.16 0.46 0.32 96.70 0.44 0.58 98.43 0.94 0.21 97.92 0.48 0.17
NUCFL (MDCA+COS) 98.80 0.52 0.17 98.18 0.35 0.25 96.17 0.41 0.42 98.14 1.03 0.21 97.83 0.52 0.18
NUCFL (DCA+L-CKA) 98.97 0.48 0.16 98.58 0.39 0.19 96.43 0.48 0.63 98.57 0.53 0.13 97.79 0.48 0.16
NUCFL (MDCA+L-CKA) 98.95 0.52 0.16 98.61 0.41 0.18 95.99 0.44 0.61 98.60 0.46 0.11 97.50 0.55 0.16
NUCFL (DCA+RBF-CKA) 99.07 0.44 0.16 98.67 0.32 0.18 96.85 0.45 0.59 98.53 0.51 0.12 97.94 0.57 0.17
NUCFL (MDCA+RBF-CKA) 98.94 0.55 0.17 98.58 0.37 0.20 96.78 0.43 0.63 98.16 0.48 0.12 97.89 0.53 0.16
@@
Table 8: Accuracy (%), calibration measures ECE (%), and SCE (%) of various federated optimization methods with different calibration methods under non-IID ($\alpha=0.5$) scenario on the MNIST dataset. Values in boldface indicate the best calibration across all methods, achieved without compromising accuracy compared to the uncalibrated model.

### v3

@@ B.2 Additional Results of Calibration on FL
Calibration Method FedAvg (McMahan et al., 2017) FedProx (Sahu et al., 2018) Scaffold (Karimireddy et al., 2019)
Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$
Uncal. 96.20 $\pm$ 0.41 0.88 $\pm$ 0.15 0.29 $\pm$ 0.08 96.17 $\pm$ 0.52 1.04 $\pm$ 0.14 0.51 $\pm$ 0.08 96.41 $\pm$ 1.02 1.89 $\pm$ 0.24 0.66 $\pm$ 0.10
Focal (Lin et al., 2017) 95.23 $\pm$ 0.52 2.71 $\pm$ 0.18 0.66 $\pm$ 0.10 97.10 $\pm$ 0.77 4.85 $\pm$ 1.10 1.11 $\pm$ 0.95 96.28 $\pm$ 1.22 2.85 $\pm$ 0.95 0.81 $\pm$ 0.44
LS (Müller et al., 2019) 96.30 $\pm$ 0.29 8.66 $\pm$ 1.88 3.31 $\pm$ 0.32 95.11 $\pm$ 0.95 7.49 $\pm$ 0.92 2.95 $\pm$ 0.24 97.00 $\pm$ 1.02 6.88 $\pm$ 1.61 2.67 $\pm$ 0.14
BS (Brier, 1950) 95.29 $\pm$ 0.98 2.74 $\pm$ 0.27 0.67 $\pm$ 0.10 95.14 $\pm$ 0.33 3.71 $\pm$ 0.10 0.76 $\pm$ 0.59 95.43 $\pm$ 0.61 2.18 $\pm$ 0.43 0.51 $\pm$ 0.12
MMCE (Kumar et al., 2018) 97.61 $\pm$ 0.62 4.88 $\pm$ 0.20 1.14 $\pm$ 0.11 97.23 $\pm$ 0.53 5.60 $\pm$ 0.33 1.18 $\pm$ 0.11 94.77 $\pm$ 1.22 12.19 $\pm$ 0.19 3.79 $\pm$ 0.10
FLSD (Mukhoti et al., 2020) 95.23 $\pm$ 0.33 8.64 $\pm$ 0.17 3.30 $\pm$ 0.09 95.44 $\pm$ 0.52 9.02 $\pm$ 0.23 3.42 $\pm$ 0.11 95.52 $\pm$ 0.24 10.60 $\pm$ 0.25 2.08 $\pm$ 0.14
MbLS (Liu et al., 2021) 95.50 $\pm$ 1.02 3.01 $\pm$ 1.30 1.10 $\pm$ 0.29 95.61 $\pm$ 0.66 4.11 $\pm$ 0.44 1.10 $\pm$ 0.25 96.31 $\pm$ 0.85 3.05 $\pm$ 0.47 0.77 $\pm$ 0.11
DCA (Liang et al., 2020) 96.43 $\pm$ 0.45 0.52 $\pm$ 0.08 0.20 $\pm$ 0.06 97.02 $\pm$ 0.50 0.68 $\pm$ 0.10 0.22 $\pm$ 0.07 97.71 $\pm$ 1.02 0.70 $\pm$ 0.11 0.23 $\pm$ 0.07
NUCFL (DCA+COS) 97.10 $\pm$ 0.52 0.49 $\pm$ 0.08 0.18 $\pm$ 0.05 96.94 $\pm$ 0.77 0.59 $\pm$ 0.09 0.19 $\pm$ 0.06 97.00 $\pm$ 1.22 0.69 $\pm$ 0.10 0.23 $\pm$ 0.07
NUCFL (DCA+L-CKA) 96.90 $\pm$ 0.41 0.44 $\pm$ 0.21 0.16 $\pm$ 0.10 97.07 $\pm$ 0.52 0.45 $\pm$ 0.07 0.15 $\pm$ 0.11 96.84 $\pm$ 1.02 0.56 $\pm$ 0.09 0.20 $\pm$ 0.06
NUCFL (DCA+RBF-CKA) 96.01 $\pm$ 0.52 0.44 $\pm$ 0.07 0.17 $\pm$ 0.05 96.91 $\pm$ 0.77 0.43 $\pm$ 0.20 0.16 $\pm$ 0.13 96.93 $\pm$ 1.22 0.53 $\pm$ 0.08 0.21 $\pm$ 0.06
MDCA (Hebbalaguppe et al., 2022b) 96.87 $\pm$ 0.41 0.61 $\pm$ 0.09 0.24 $\pm$ 0.07 97.04 $\pm$ 0.52 0.85 $\pm$ 0.13 0.30 $\pm$ 0.08 97.04 $\pm$ 1.02 0.92 $\pm$ 0.14 0.33 $\pm$ 0.09
NUCFL (MDCA+COS) 97.11 $\pm$ 0.64 0.54 $\pm$ 0.10 0.20 $\pm$ 0.09 96.95 $\pm$ 0.22 0.56 $\pm$ 0.10 0.19 $\pm$ 0.09 96.97 $\pm$ 1.00 0.61 $\pm$ 0.83 0.22 $\pm$ 0.20
NUCFL (MDCA+L-CKA) 96.89 $\pm$ 0.85 0.52 $\pm$ 0.27 0.20 $\pm$ 0.09 97.02 $\pm$ 0.25 0.55 $\pm$ 0.49 0.20 $\pm$ 0.10 97.11 $\pm$ 1.04 0.55 $\pm$ 0.08 0.20 $\pm$ 0.02
NUCFL (MDCA+RBF-CKA) 97.01 $\pm$ 0.95 0.55 $\pm$ 0.11 0.21 $\pm$ 0.06 96.97 $\pm$ 0.57 0.64 $\pm$ 0.14 0.23 $\pm$ 0.06 97.54 $\pm$ 0.83 0.70 $\pm$ 0.25 0.24 $\pm$ 0.07
Calibration Method FedDyn (Acar et al., 2021) FedNova (Wang et al., 2020b)
Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$
Uncal. 97.23 $\pm$ 1.22 1.40 $\pm$ 0.69 0.55 $\pm$ 0.13 96.42 $\pm$ 1.04 0.89 $\pm$ 0.49 0.28 $\pm$ 0.11
Focal (Lin et al., 2017) 97.61 $\pm$ 0.54 4.37 $\pm$ 0.85 1.09 $\pm$ 0.10 95.23 $\pm$ 0.92 3.82 $\pm$ 0.94 0.80 $\pm$ 0.09
LS (Müller et al., 2019) 95.11 $\pm$ 0.59 5.02 $\pm$ 1.02 2.50 $\pm$ 0.20 96.61 $\pm$ 0.99 10.42 $\pm$ 1.31 3.62 $\pm$ 0.81
BS (Brier, 1950) 92.63 $\pm$ 1.04 8.49 $\pm$ 0.44 0.88 $\pm$ 0.09 96.10 $\pm$ 1.00 3.13 $\pm$ 0.17 0.70 $\pm$ 0.07
MMCE (Kumar et al., 2018) 97.10 $\pm$ 0.79 3.04 $\pm$ 0.26 0.79 $\pm$ 0.07 96.10 $\pm$ 0.82 5.44 $\pm$ 0.41 1.12 $\pm$ 0.15
FLSD (Mukhoti et al., 2020) 95.41 $\pm$ 0.39 3.29 $\pm$ 0.31 0.72 $\pm$ 0.11 95.44 $\pm$ 0.41 7.63 $\pm$ 0.52 2.81 $\pm$ 0.15
MbLS (Liu et al., 2021) 95.22 $\pm$ 0.92 3.09 $\pm$ 0.33 0.79 $\pm$ 0.10 97.07 $\pm$ 0.85 4.17 $\pm$ 0.42 1.06 $\pm$ 0.17
DCA (Liang et al., 2020) 97.39 $\pm$ 1.22 0.57 $\pm$ 0.09 0.30 $\pm$ 0.08 97.11 $\pm$ 1.04 0.64 $\pm$ 0.10 0.27 $\pm$ 0.07
NUCFL (DCA+COS) 97.61 $\pm$ 0.54 0.54 $\pm$ 0.08 0.25 $\pm$ 0.07 97.42 $\pm$ 0.92 0.57 $\pm$ 0.09 0.29 $\pm$ 0.08
NUCFL (DCA+L-CKA) 97.53 $\pm$ 1.22 0.46 $\pm$ 0.07 0.17 $\pm$ 0.09 97.18 $\pm$ 1.04 0.26 $\pm$ 0.13 0.16 $\pm$ 0.05
NUCFL (DCA+RBF-CKA) 96.75 $\pm$ 0.54 0.44 $\pm$ 0.07 0.17 $\pm$ 0.09 97.63 $\pm$ 0.92 0.29 $\pm$ 0.05 0.17 $\pm$ 0.05
MDCA (Hebbalaguppe et al., 2022b) 97.41 $\pm$ 1.22 0.65 $\pm$ 0.10 0.28 $\pm$ 0.08 97.04 $\pm$ 1.04 0.75 $\pm$ 0.11 0.28 $\pm$ 0.08
NUCFL (MDCA+COS) 96.93 $\pm$ 0.25 0.51 $\pm$ 1.00 0.25 $\pm$ 0.09 96.99 $\pm$ 1.11 0.62 $\pm$ 0.22 0.26 $\pm$ 0.10
NUCFL (MDCA+L-CKA) 97.09 $\pm$ 0.25 0.50 $\pm$ 0.33 0.26 $\pm$ 0.06 97.14 $\pm$ 1.00 0.54 $\pm$ 0.38 0.24 $\pm$ 0.09
NUCFL (MDCA+RBF-CKA) 97.86 $\pm$ 0.61 0.55 $\pm$ 0.09 0.26 $\pm$ 0.06 97.58 $\pm$ 0.52 0.60 $\pm$ 0.23 0.26 $\pm$ 0.10
@@
Table 7: Accuracy (%), calibration measures ECE (%), and SCE (%) of various federated optimization methods with different calibration methods under IID scenario on the MNIST dataset.
@@
Calibration Method FedAvg (McMahan et al., 2017) FedProx (Sahu et al., 2018) Scaffold (Karimireddy et al., 2019)
Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$
Uncal. 95.02 $\pm$ 0.41 1.89 $\pm$ 0.15 0.65 $\pm$ 0.08 94.55 $\pm$ 0.52 2.04 $\pm$ 0.14 0.65 $\pm$ 0.08 94.15 $\pm$ 1.02 4.54 $\pm$ 0.24 1.11 $\pm$ 0.10
Focal (Lin et al., 2017) 94.38 $\pm$ 0.52 4.46 $\pm$ 0.18 1.25 $\pm$ 0.10 93.44 $\pm$ 0.77 4.17 $\pm$ 1.10 1.20 $\pm$ 0.95 93.51 $\pm$ 1.22 7.65 $\pm$ 0.95 1.52 $\pm$ 0.44
LS (Müller et al., 2019) 94.53 $\pm$ 0.29 8.00 $\pm$ 1.88 3.30 $\pm$ 0.32 95.15 $\pm$ 0.95 3.91 $\pm$ 0.92 1.20 $\pm$ 0.24 95.25 $\pm$ 1.02 5.66 $\pm$ 1.31 4.50 $\pm$ 0.14
BS (Brier, 1950) 93.26 $\pm$ 0.98 4.02 $\pm$ 0.27 1.17 $\pm$ 0.10 92.58 $\pm$ 0.33 4.78 $\pm$ 0.10 1.26 $\pm$ 0.59 92.77 $\pm$ 0.61 4.44 $\pm$ 0.43 1.10 $\pm$ 0.12
MMCE (Kumar et al., 2018) 95.43 $\pm$ 0.62 7.82 $\pm$ 0.20 2.61 $\pm$ 0.11 93.74 $\pm$ 0.53 6.56 $\pm$ 0.33 1.95 $\pm$ 0.11 88.48 $\pm$ 1.22 7.88 $\pm$ 0.19 1.58 $\pm$ 0.10
FLSD (Mukhoti et al., 2020) 93.95 $\pm$ 0.33 8.46 $\pm$ 0.17 3.48 $\pm$ 0.09 92.29 $\pm$ 0.52 8.81 $\pm$ 0.23 3.88 $\pm$ 0.11 92.41 $\pm$ 0.24 7.38 $\pm$ 0.25 1.53 $\pm$ 0.14
MbLS (Liu et al., 2021) 94.22 $\pm$ 1.02 5.99 $\pm$ 1.30 1.83 $\pm$ 0.29 93.46 $\pm$ 0.66 4.84 $\pm$ 0.44 1.25 $\pm$ 0.25 95.55 $\pm$ 0.85 6.20 $\pm$ 0.47 2.41 $\pm$ 0.11
DCA (Liang et al., 2020) 95.94 $\pm$ 0.40 1.09 $\pm$ 0.10 0.49 $\pm$ 0.05 95.04 $\pm$ 0.50 1.87 $\pm$ 0.12 0.59 $\pm$ 0.06 95.42 $\pm$ 0.60 3.75 $\pm$ 0.08 1.08 $\pm$ 0.07
NUCFL (DCA+COS) 96.16 $\pm$ 0.35 0.81 $\pm$ 0.05 0.34 $\pm$ 0.03 95.16 $\pm$ 0.45 1.66 $\pm$ 0.04 0.56 $\pm$ 0.05 95.70 $\pm$ 0.55 2.91 $\pm$ 0.06 0.85 $\pm$ 0.08
NUCFL (DCA+L-CKA) 95.97 $\pm$ 0.35 0.78 $\pm$ 0.04 0.31 $\pm$ 0.02 95.58 $\pm$ 0.45 1.69 $\pm$ 0.05 0.56 $\pm$ 0.03 95.43 $\pm$ 0.55 2.78 $\pm$ 0.07 0.81 $\pm$ 0.08
NUCFL (DCA+RBF-CKA) 96.07 $\pm$ 0.37 0.74 $\pm$ 0.04 0.30 $\pm$ 0.02 95.67 $\pm$ 0.45 1.66 $\pm$ 0.04 0.55 $\pm$ 0.03 95.85 $\pm$ 0.55 3.05 $\pm$ 0.06 0.87 $\pm$ 0.08
MDCA (Hebbalaguppe et al., 2022b) 95.92 $\pm$ 0.41 1.38 $\pm$ 0.12 0.55 $\pm$ 0.06 94.69 $\pm$ 0.52 1.93 $\pm$ 0.14 0.62 $\pm$ 0.08 94.82 $\pm$ 1.02 3.26 $\pm$ 0.15 1.04 $\pm$ 0.09
NUCFL (MDCA+COS) 95.80 $\pm$ 0.40 0.88 $\pm$ 0.05 0.42 $\pm$ 0.03 95.18 $\pm$ 0.50 1.69 $\pm$ 0.04 0.57 $\pm$ 0.04 95.17 $\pm$ 0.60 3.01 $\pm$ 0.06 0.87 $\pm$ 0.07
NUCFL (MDCA+L-CKA) 95.95 $\pm$ 0.47 0.82 $\pm$ 0.09 0.43 $\pm$ 0.02 95.61 $\pm$ 0.50 1.64 $\pm$ 0.05 0.55 $\pm$ 0.03 94.99 $\pm$ 0.54 2.74 $\pm$ 0.06 0.80 $\pm$ 0.08
NUCFL (MDCA+RBF-CKA) 95.94 $\pm$ 0.41 0.85 $\pm$ 0.05 0.42 $\pm$ 0.03 94.58 $\pm$ 0.58 1.67 $\pm$ 0.10 0.58 $\pm$ 0.10 95.08 $\pm$ 0.75 2.83 $\pm$ 0.06 0.68 $\pm$ 0.08
Calibration Method FedDyn (Acar et al., 2021) FedNova (Wang et al., 2020b)
Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$
Uncal. 96.87 $\pm$ 1.22 1.78 $\pm$ 0.69 0.61 $\pm$ 0.13 95.05 $\pm$ 1.04 1.54 $\pm$ 0.49 0.54 $\pm$ 0.11
Focal (Lin et al., 2017) 96.78 $\pm$ 0.54 2.21 $\pm$ 0.85 0.77 $\pm$ 0.10 94.56 $\pm$ 0.92 4.38 $\pm$ 0.94 1.10 $\pm$ 0.09
LS (Müller et al., 2019) 97.80 $\pm$ 0.59 5.65 $\pm$ 1.02 2.96 $\pm$ 0.20 96.48 $\pm$ 0.99 3.41 $\pm$ 1.31 1.94 $\pm$ 0.81
BS (Brier, 1950) 93.68 $\pm$ 1.04 5.81 $\pm$ 0.44 2.93 $\pm$ 0.09 92.96 $\pm$ 1.00 3.65 $\pm$ 0.17 1.98 $\pm$ 0.07
MMCE (Kumar et al., 2018) 96.48 $\pm$ 0.79 2.08 $\pm$ 0.26 0.74 $\pm$ 0.07 96.30 $\pm$ 0.82 6.75 $\pm$ 0.41 2.57 $\pm$ 0.15
FLSD (Mukhoti et al., 2020) 94.84 $\pm$ 0.39 2.49 $\pm$ 0.31 0.74 $\pm$ 0.11 93.93 $\pm$ 0.41 5.57 $\pm$ 0.52 2.45 $\pm$ 0.15
MbLS (Liu et al., 2021) 95.71 $\pm$ 0.92 3.95 $\pm$ 0.33 1.06 $\pm$ 0.10 94.24 $\pm$ 0.85 5.67 $\pm$ 0.42 2.82 $\pm$ 0.17
DCA (Liang et al., 2020) 96.47 $\pm$ 0.70 1.44 $\pm$ 0.09 0.47 $\pm$ 0.03 95.13 $\pm$ 0.50 0.94 $\pm$ 0.08 0.48 $\pm$ 0.05
NUCFL (DCA+COS) 96.93 $\pm$ 0.65 1.24 $\pm$ 0.07 0.43 $\pm$ 0.04 96.01 $\pm$ 0.55 0.78 $\pm$ 0.05 0.45 $\pm$ 0.03
NUCFL (DCA+L-CKA) 97.07 $\pm$ 0.65 1.23 $\pm$ 0.06 0.43 $\pm$ 0.02 96.05 $\pm$ 0.55 0.77 $\pm$ 0.05 0.44 $\pm$ 0.03
NUCFL (DCA+RBF-CKA) 97.03 $\pm$ 0.65 1.31 $\pm$ 0.05 0.47 $\pm$ 0.02 96.00 $\pm$ 0.55 0.85 $\pm$ 0.06 0.45 $\pm$ 0.03
MDCA (Hebbalaguppe et al., 2022b) 97.00 $\pm$ 1.22 1.40 $\pm$ 0.16 0.47 $\pm$ 0.04 95.10 $\pm$ 1.04 1.30 $\pm$ 0.18 0.50 $\pm$ 0.07
NUCFL (MDCA+COS) 96.64 $\pm$ 0.70 1.33 $\pm$ 0.08 0.45 $\pm$ 0.04 96.02 $\pm$ 0.60 0.82 $\pm$ 0.06 0.47 $\pm$ 0.03
NUCFL (MDCA+L-CKA) 97.10 $\pm$ 0.77 1.30 $\pm$ 0.22 0.44 $\pm$ 0.08 96.50 $\pm$ 0.65 0.80 $\pm$ 0.10 0.45 $\pm$ 0.09
NUCFL (MDCA+RBF-CKA) 96.66 $\pm$ 0.70 1.48 $\pm$ 0.05 0.49 $\pm$ 0.02 96.19 $\pm$ 0.56 0.85 $\pm$ 0.11 0.47 $\pm$ 0.10
@@
Table 8: Accuracy (%), calibration measures ECE (%), and SCE (%) of various federated optimization methods with different calibration methods under non-IID ($\alpha=0.5$) scenario on the MNIST dataset. Underlined values indicate the best calibration across all methods.